OpenStudy (anonymous):

Can someone explain the technique if completing the square to transform the quadratic equation into the form (x+c)^2=a??

OpenStudy (amistre64):

a complete square is a perfect square; can you expand out the x+c part?

Parth (parthkohli):

Basically, if you have something like \(ax^2 + bx + c=0\) and \(ax^2 + bx + c\) is not a perfect square, you add \(\left(b/2\right)^2\) to both sides to make it so.

Parth (parthkohli):

If that is what you ask

OpenStudy (anonymous):

Thank you

Parth (parthkohli):

Do you know WHY we do that?

OpenStudy (anonymous):

I get the ax^2+bx+c=0 but I'm not so sure about the (b/2) thing

Parth (parthkohli):

It's pretty simple, in fact! Hint:\[(x + a)^2 = x^2 + (2a)x + a^2\]What do you do to get \((a)^2\) from \((2a)x\)? You simply take the coefficient of \(x\), divide by \(2\), and square it.

Parth (parthkohli):

Coefficient of \(x\): \({2a}\). Divide both sides by \(2\): \(a\). Square: \(a^2\).

Parth (parthkohli):

That will get you the constant term.

OpenStudy (anonymous):

Okay. That part makes sense but I'm still not sure about the (b/2)^2

Parth (parthkohli):

What does \(b\) mean? The coefficient of \(x\). What is the coefficient of \(x\)? \(2a\)...

OpenStudy (anonymous):

\[6x^2 +36x+18=0\] So in this case, it would be 36?

Parth (parthkohli):

Yes, right there!

Parth (parthkohli):

Divide \(36\) by \(2\), square...

Parth (parthkohli):

Before doing that, take \(18\) to the right side of the equation to make life easier.

OpenStudy (anonymous):

So put the equation as \[6x^2+36=18\] then do the \[(\frac{ b }{ 2 })^2\] thing???